OpenStudy (anonymous):

knowing 2^n(A)= n(P(A)), if n(P(A))=y, then n(A)=

OpenStudy (dumbcow):

\[2^{n(A)} = y\] \[\ln (2^{n(A)}) = \ln y\] \[n(A) \ln 2 = \ln y\] \[n(A) = \frac{\ln y}{\ln 2}\]